Two events A and B are such that 𝑃(𝐴) = 0.45, 𝑃(𝐡) = 0.22 π‘Žπ‘›π‘‘ 𝑃(𝐴 βˆͺ 𝐡) = 0.53.

5.1.1 Find 𝑃(𝐴̅βˆͺ 𝐡̅)

5.1.2 Are the events A and B independent?Β 


1
Expert's answer
2021-09-21T12:31:28-0400

We can present the union of events as: "A\cup B=(A\backslash(A\cap B))\cup(A\cap B)\cup(B\backslash(A\cap B))". Thus, we get: "P(A\cup B)=P(A\backslash(A\cap B))+P((A\cap B))+P(B\backslash(A\cap B))=P(A)+P(B\backslash(A\cap B))".

From the latter we find: "P(A\cup B)-P(A)=P(B\backslash(A\cap B))".

From the latter we find that: "0.53-0.45=P(B)-P(A\cap B)" . We substitute "P(B)=0.22" and get: "P(A\cap B)=0.22-0.08=0.14".

Denote the whole probability space by "X". Then, "\bar{A}=X\backslash A" and "\bar{B}=X\backslash B". Thus, we get: "\bar{A}\cup{\bar{B}}=(X\backslash A)\cup(X\backslash B)=X\backslash(A\cap B)". We get:

a). "P(\bar{A}\cup{\bar{B}})=P(X\backslash(A\cap B))=P(X)-P(A\cap B)=1-0.14=0.86".

b). For independent events one has: "P(A\cap B)=P(A)P(B)". "P(A)P(B)=0.099". It follows from the formulation of the problem. As we can see "P(A)P(B)\neq P(A\cap B)". Thus, the events are not independent.